75 kg ball player moving at a velocity of 7.8m/s hits a player moving in the opposite direction who has a mass of 90 kg. If they collide and stop to have a speed of zero how fast is the other player moving originally?



1
Expert's answer
2021-05-10T08:28:02-0400

We can find the initial velocity of the other player from the law of conservation of momentum:


"p_i=p_f,""m_1v_{1i}+m_2v_{2i}=0,""v_{2i}=-\dfrac{m_1v_{1i}}{m_2}=-\dfrac{75\ kg\cdot7.8\ \dfrac{m}{s}}{90\ kg}=-6.5\ \dfrac{m}{s}."

The sign minus means that the other player moves in the opposite direction before the collision.

Answer:

"v_{2i}=6.5\ \dfrac{m}{s}", in the opposite direction.